OpenStudy (anonymous):

rationalize 9/5-√7 multiply by the conjugate right?

OpenStudy (angela210793):

yes...:):)

OpenStudy (anonymous):

could you do it because I'm getting 45+9√7 :P

OpenStudy (anonymous):

is this \[\frac{9}{5-\sqrt{7}}\]?

OpenStudy (anonymous):

sha, sate.

OpenStudy (angela210793):

9(5+sqr7)/18=(45+9sqr5)/18

OpenStudy (anonymous):

18s. random 18s.

OpenStudy (anonymous):

if so then \[\frac{9}{5-\sqrt{7}}\times \frac{5+\sqrt{7}}{5+\sqrt{7}}\] \[=\frac{9(5+\sqrt{7}}{25-7}=\frac{9(5-\sqrt{7})}{18}\]

OpenStudy (anonymous):

totally get it. thanks broskis and shebroski.

OpenStudy (angela210793):

broskis and wht? wht is tht? O.o

OpenStudy (anonymous):

wait is it (45+√5?)/18?

OpenStudy (anonymous):

ski is a slang suffix.. so bro, and ski. broski.

OpenStudy (angela210793):

just leave it in the way...(45+√5?)/18?

OpenStudy (anonymous):

people are into adding weird suffixi to the word bro. BROSAMABINLADEN. Brosuarusrex. Broseph. Brotato.

OpenStudy (anonymous):

well shouldn't it be (45+√7)/18?

OpenStudy (angela210793):

yea it should be in tht way sorry forgot 7..it must be (45+√7)/18

OpenStudy (anonymous):

thank you!

OpenStudy (angela210793):

yw